OpenStudy (anonymous):

When solving the system x+2y=9 and x-y=6, which expression could be substituted for x in the second equation?

OpenStudy (anonymous):

A. 9-x B. y=9/2 C. 9-2y D. 1/2y+9

OpenStudy (anonymous):

Isolate x in the first equation, then substitute it into the second. x + 2 y = 9 x = 9 - 2y Plug that into the second equation for x and solve.